Pressure shifts in the ν_{3}-band of NO_{2} induced by collisions with noble gases, N_{2}, H_{2}, D_{2}, O_{2}, CO_{2}, and SO_{2}

\underline{A. KISSEL}, Optisches Institut der Technischen Universit t Berlin, Sekr. PN 0-1, Hardenbergstr. 36, 10623 Berlin, Germany; S. BOUAZZA, permanent address: Département de Physique, Université de Champagne-Ardenne Reims, France; B. SUMPF*, H.-D. KRONFELDT, Optisches Institut der Technischen Universit t Berlin, Sekr. PN 0-1, Hardenbergstr. 36, 10623 Berlin, Germany.

Very high resolution (0.0005 cm^{−1}) infrared spectra of NO_{2} have been recorded with a pulse driven diode laser spectrometer. The 1600 cm^{−1} region corresponding to the ν_{3}-band has been analysed. 34 lines (6 ≤ N′′ ≤ 38, 0 ≤ K′′_{a} ≤ 11), with NO_{2} pressure varying from 0.1 to 0.5 Torr and foreign gas (Xe, Kr, Ar,...) from 100 to 220 Torr were recorded, using a Herriot cell of 4.7 m pathlength due the weakness of lines. The results clearly shows (within experimental accuracy) a notable variation of shift with respect to N′′ quantum number, the shifts being larger in negativ magnitude for big N′′. One can also note that the values of shifts are not systematically lower for molecules with higher broadening coefficients as it is for H_{2}O^{a}. The similarities and differences between H_{2}O and NO_{2} will be discussed.

*B.S. acknowledges a grant from the Deutsche Forschungsgemeinschaft

aB.Grossmann and E.Browell, J.Mol. Spectrosc. 138, 562 (1989)